# BWTP model for hadrons $p_{\rm T}$-spectra and elliptic flow $v_{2}$ in AA collisions at the LHC

### v1

@@ II Description of BWTPM
Figure 1: Fit of the midrapidity charged-particle multiplicity density data measured for various AA collisions at different centralities $x_{c}$ and energies $\sqrt{s_{\mathrm{NN}}}$ by the ALICE [17, 18, 19, 20], PHOBOS [21] and PHENIX [22] experiments. Dashed line shows our prediction (multiplied by 0.5 for better visibility) for the OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV. Ratio data/fit demonstrates the quality of the fits.
@@
To our knowledge, there is no similar to Eq. (13) generic parameterization in the literature. The existing parameterizations of $M_{\rm ch}$ versus $x_{c}$ (or versus the number of participant nucleons $N_{\mathrm{part}}$) are valid only for a particular AA collision system [21, 29]. Of course, there are many sophisticated models describing the $M_{\rm ch}$ data, most of which use the hydrodynamical approach for particle production in AA collisions (see, e.g., Ref. [30]). However, such models are much more complex with respect to our simple parameterization and their accuracy is worse.
@@
Figure 2: Fit of the midrapidity charged-particle multiplicity density in inelastic pp collisions by Eq. (14) as a function of $\sqrt{s}$. Higher energy six data-points are from Refs. [24, 25, 26] and the remaining data-points are from Refs. [21, 27].
IV parameterization of BWTPM
@@
Here we present parameterizations for the model parameters versus the $M_{\rm ch}$ (given by Eq. (13)), A, $\sqrt{s_{\mathrm{NN}}}$, and the particle mass and $p_{\textrm{T}}$. The parameterizations were chosen as simple as possible by examining the behavior of the model parameters versus the collision system, centrality, and energy, and using the results of Ref. [1] and other studies. For normalization parameters $V,V_{1}$, and $V_{2}$, to describe their strong increase with centrality [1] and taking into account the experimental fact that the hadron $p_{\textrm{T}}$-integrated yields increase linearly with $M_{\rm ch}$ at $M_{\rm ch}>100$ [31], the following parameterizations are assumed
@@
Here and throughout the present paper the parameters $a_{i},i=1,2,3,$... are fitting constants. Parameter $f_{r}$ is nonzero for ${\mathrm{K}}^{*0}$and $\rho^{0}$ only, and accounts for the suppression of the yields of these short-lived resonances at low $p_{\textrm{T}}$ due to the rescatterings of their decay products in the hadronic medium. The suppression is stronger for more central collisions and scales with $M_{\rm ch}^{1/3}$ [11].
The decrease of parameters $T$ and $k$ with increase of centrality and of parameters $e_{2}$ and $n$ with increase of centrality and $\sqrt{s_{\mathrm{NN}}}$ [1] can be parameterized as follows
@@
The factor $f_{S}$ is equal unity for ordinary hadrons and increases with decreasing centrality (or $M_{\rm ch}$) for other hadrons. It is greater for hadrons containing larger number of strange and charm (anti)quarks, which we denote by $n_{s}$ and $n_{c}$, respectively. For example, $n_{s}=1$ for K and $\eta$  mesons, and $n_{s}=2$ for $\phi$ meson. We have found a simple parameterization for $f_{S}$ in AA collisions
@@
which describes well the enhancement features of all considered hadrons containing at least one strange or charm (anti)quark ($n_{s}=n_{c}=0$ for ordinary hadrons). The term proportional to the hadron baryon number $B$ is introduced to better describe the production of hyperons. We assume the same parameter $a_{21}$ for strange and charmed baryons since the used data for the latter case are very limited. For the hadrons containing bottom quarks, not considered in the present study, the Eq. (20) can be modified accordingly. Note that $f_{S}$ tends to unity at $p_{\textrm{T}}>a_{22}$. Hence, the ratios of different particle yields in our model are independent of centrality at high $p_{\textrm{T}}$, as in the data [11].
It is known that the $p_{\textrm{T}}$-averaged $v_{2}$ increases when going from central to peripheral collisions, and then decreases for very peripheral collisions [11, 16]. Also, the eccentricity $\varepsilon_{2}$ of the transverse geometry of hadronic medium behaves similarly versus collision centrality. When going from the central to peripheral collisions, this geometry goes from circular to elliptical and $\varepsilon_{2}$ increases. But the geometry goes again to circular and $\varepsilon_{2}$ decreases at very peripheral collisions, where in average only one nucleon-nucleon collision accurs [10, 15]. So one can assume $v_{2}\sim\varepsilon_{2}$. This parameterization of $r_{xy}$ ensures such a centrality dependence for $\varepsilon_{2}$ (given by Eq. (7))
@@
where $x=\textrm{A}^{a_{25}}(z_{c}+y)/(1+y)$, A is the atomic mass number, $z_{c}=M_{\rm ch}/M_{\rm ch}^{0\textrm{-}5\%}$ is the normalized multiplicity over the multiplicity of the 0-5% most central collisions $M_{\rm ch}^{0\textrm{-}5\%}$ (determined by Eq. (13) at $x_{c}=0.025$) and its correction $y=(a_{31}/\sqrt{s_{\mathrm{NN}}})^{a_{32}}z_{c}$ is needed mainly for the fit of $v_{2}$ data at RHIC energies. Parameter $f_{d}$ is introduced to describe a slightly different centrality dependence of $v_{2}$ data in Xe–Xe collisions compared to all other collision systems considered in present study for which we use $f_{d}=0$. The difference is related to the non-spherical (deformed) shape of 129Xe nucleus [32].
To describe the centrality and energy dependence of the hadrons $p_{\textrm{T}}$-spectra and also of the charged-particle average transverse momentum $\langle p_{\textrm{T}}\rangle$ [33, 34], particularly its local minimum at $z_{c}\sim 1.05$ and rise at $z_{c}>1.05$ for ultracentral collisions [34], we parameterize $\rho_{0}$ as
@@
Factor $f_{uc}$ [34, 35] is close to unity at $z_{c}\!\leq\!a_{38}\!\sim\!1.1$ and decreases at $z_{c}\!>\!a_{39}$. So, it increases $\rho_{0}$ at $z_{c}\!>\!a_{39}$ and therefore increases $\langle p_{\textrm{T}}\rangle$ for ultracentral events like in Refs. [34, 35]. Parameter $a_{35}$ is used to describe the observed slightly different centrality dependence of $\langle p_{\textrm{T}}\rangle$ in Xe–Xe and Pb–Pb collisions which is due to the non-sphericity of 129Xe nuclei [33].
So, the parameterizations in Eqs. (16)–(IV) allow us to determine the aforementioned in Sec. II more than 600 independent values for $V,V_{1},V_{2},T,k,e_{2},n,f_{S},r_{xy},$ and $\rho_{0}$ by means of about 40 constants.
We further parameterize also other ingredients of BWTPM. The energy dependence of parameter $f_{hyp}$ from Eqs. (11) and (12), and of the global normalization $f_{N}$ are given as
@@
where $a_{N}=1$ is assumed for Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}=$ 5.02 TeV and for other used data-sets the values of $a_{N}$ are defined relative to this one. As was mentioned in Sec. II, we assume that the parameter $\rho_{2}$ as well as the functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ are proportional to $\varepsilon_{2}$ in order to ensure the relation $v_{2}\sim\varepsilon_{2}$ [13, 14, 15, 16]. A weak energy dependence of $\rho_{2}$ is parameterized as
$\rho_{2}=\varepsilon_{2}a_{44}/\sqrt{s_{\mathrm{NN}}}^{a_{45}},$ (24)
which becomes significant at RHIC energies.
Functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ are responsible mainly for the description of the high-$p_{\textrm{T}}$ behavior of elliptic flow $v_{2}$. For their dependence on $p_{\textrm{T}}$, $\sqrt{s_{\mathrm{NN}}}$, and $M_{\rm ch}$ we use
@@
parameterization for mesons is inaccurate for $\mathrm{J}/\psi$.
@@
where $m_{\mathrm{p}}$ is the proton mass and the step function $H$ is defined as $H(x)=0(1)$ for $x\leq(>)~0$. So $c_{2}=1$ for pions and (anti)proton. Larger $a_{59}$ with respect to $a_{58}$ is nedded to describe the baryon yields. Note that the different parameterizations for mesons and baryons in Eqs.  (26) and (27) indicate the difference in production mechanisms of these hadrons containing two and three (anti)quarks, respectively.
@@
where parameter $\tilde{e}_{1}$ depends on the particle species. Parameter $f_{\mathrm{p}}$ is nonzero for (anti)proton only to better describe its $p_{\textrm{T}}$-spectra at RHIC energies. Additional normalization parameter $f_{\rm D}$ is used for $\mathrm{D}^{\pm}$ mesons to describe their lower yield with respect to ${\mathrm{D}}^{0}$ meson.
@@ V Fitted data and parameters
Thus, we presented in previous Sections all the ingredients of the BWTPM. To determine the values of its free parameters, a global simultaneous fit is performed of the combined data on various hadrons midrapidity $p_{\textrm{T}}$-spectra, $\langle p_{\textrm{T}}\rangle$, and $v_{2}$ for different centrality classes in Pb–Pb collisions at energies $\sqrt{s_{\mathrm{NN}}}=2.76$ and 5.02 TeV [18, 33, 34, 36, 37, 40, 41, 42, 38, 39, 43, 44, 45, 48, 49, 46, 47, 50, 51, 54, 55, 52, 53, 56, 57, 58, 59, 60, 61] and in Xe–Xe collisions at 5.44 TeV [32, 62, 63, 64]. To check the validity of BWTPM for other colliding nuclei and at much lower energies, we included in this fit also the $p_{\textrm{T}}$-spectra and $v_{2}$ for $\pi^{\pm}$, $\mathrm{K}^{\pm}$ and $\mathrm{p}$($\mathrm{\overline{p}}$) measured in Au–Au collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV [65, 66], and the $p_{\textrm{T}}$-spectra for $\pi^{\pm}$ and $\mathrm{K}^{\pm}$ measured in Au–Au collisions at 39 and 27 GeV [67], in UU collisions at 193 GeV [68] and in Cu–Cu collisions at 200 GeV [69]. In the last four cases the proton data not included since they are not corrected for the contribution from weak decays of hyperons. The resulting ratio $\chi^{2}/NDF$ for this global fit (in the ROOT framework [28]) is 6738/13408. Values for the fitting parameters are given in Tables I–III. Parameters $f_{BW},f_{1}$ and $f_{2}$ for $\pi^{\pm}$ and $\pi^{0}$ are fixed equal to unity. Note that in the fits we use the data of prompt $\rm D$} and $\mathrm{J}/\psi$ mesons and $\Lambda_{c}$ baryons, not including contributions from decays of heavier hadrons, containing $b$-quarks. To describe the $p_{\textrm{T}}$-spectra in Xe–Xe collisions [62, 63], we scale the right-hand side of Eq. (12) by a factor of 1.07 to account for a small normalization difference between the measured $p_{\textrm{T}}$-spectra of identified [62] and unidentified charged particles [63]. Note also that parameters $a_{24}$$a_{32}$ and $a_{44}$$a_{57}$ are resposible mainly for the elliptic flow. Only some of them entering in the Eq. (24) influence the azimuthally integrated $p_{\textrm{T}}$-spectra due to the $\rho_{2}$ contribution into the radial flow velocity Eq. (5). This contribution is small and decreases with increasing energy, according to Eq. (24). So, its influence on $p_{\textrm{T}}$-spectra is negligible at high energies and becomes sizable (at most about several % at high $p_{\textrm{T}}$) only for Au–Au collisions at low energies $\sqrt{s_{\mathrm{NN}}}=27$ and 39 GeV.
@@ VI.1 Hadrons $p_{\textrm{T}}$-spectra, integrated yields, and $\langle p_{\textrm{T}}\rangle$
First, we present several plots to illustrate the results and quality of the $p_{\textrm{T}}$-spectra fits. Figure 3 displays the fits of different data sets with highest $p_{\textrm{T}}$ reach in central Pb–Pb collisions. To demonstrate the quality of the fits, the data points have been divided by the corresponding fit function values, and these ratios are also plotted in the bottom panel. Generally, the quality is good within the data uncertainties. Figures 4, 5, and 6 display our main fits of the centrality dependent spectra for most abundantly produced pions, kaons, and protons. The fit quality is almost always very good. This is true also for the fits of $p_{\textrm{T}}$-spectra of all considered other particle species. Our predictions for the 0-5% most central OO collisions are shown in Fig. 5. Note that Figs. 3 and 4
Figure 3: Fit of $p_{\textrm{T}}$-spectra with highest $p_{\textrm{T}}$ reach in Pb–Pb collisions for charged particles at $\sqrt{s_{\mathrm{NN}}}$ ==2.76 TeV [37, 40] and 5.02 TeV [37, 41] and for ${\mathrm{D}}^{0}$  at $\sqrt{s_{\mathrm{NN}}}$ ==5.02 TeV [42, 47].
Figure 4: Fit of pion, kaon, and proton $p_{\textrm{T}}$-spectra at $|y|<$ 0.5 for different centrality classes (centrality intervals) in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV [18]. The data points and fitting curves in the top panels are scaled by the numbers given in the parentheses for a better visibility.
Figure 5: Similar to the Fig. 4 but for Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [62]. In addition, the dashed lines in top panels show our predictions (scaled down by 0.0002 for better visibility) for 0-5 % centrality OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
@@
repeat the similar ones of Ref. [1] to show that the present version of BWTPM with much less number of fitting parameters gives similar fit quality. Also, we have checked that the relative contributions of the BW, Ts, and Po components of present model into the $p_{\textrm{T}}$-spectra are similar to the ones given in Ref. [1]. As shown in Table I, the contribution of the Ts-component is nonzero only for pions, kaons, and protons, which have significant feedback from the resonance decays.
@@
Then, we show in Fig. 7 a good agreement of the model calculations (using the parameter values given in Sec. V) for the ratios of $p_{\textrm{T}}$-spectra of different hadrons with the ALICE data obtained in central and peripheral Pb–Pb collisions. These ratios, like the other ones given in Ref. [1], demonstrate more clearly the effect of hadron production mechanisms on the $p_{\textrm{T}}$-spectra. The peaks at $p_{\textrm{T}}$ $\sim$ 2–7 GeV/c are related to the radial flow, which is stronger for heavier particles and more central collisions. The ratios with the short-lived ${\mathrm{K}}^{*0}$and $\rho^{0}$ resonances show the suppression of their yields for low $p_{\textrm{T}}$ and central collisions, according to Eq. (16). To demonstrate the collision system and energy dependence of particle ratios, we show in the left panels of Fig. 7 the model calculations for 0-10% centrality Au–Au (dashed lines) and Cu–Cu (dotted lines) collisions at $\sqrt{s_{\mathrm{NN}}}$ == 200 GeV. Both dependencies are not strong and appear at moderate $p_{\textrm{T}}$. Note that the hadron $p_{\textrm{T}}$-spectra in Eq. (II) have a power-law behavior $\propto p_{\textrm{T}}^{2-n}$ at high $p_{\textrm{T}}$. Hence, the ratios for all hadrons, which have $h_{n}=1$ in Table I, reach a plateau (independent of the collision system, centrality, and $\sqrt{s_{\mathrm{NN}}}$) at high $p_{\textrm{T}}$ where parameter $n$ in Eq. (19) depends only on $\sqrt{s_{\mathrm{NN}}}$. The plateau is simply defined by the ratio of factors $(2J+1)f_{2}$ with $f_{2}$ given in Table I. It is true also for the ratios between hyperons which have the same $h_{n}$. For example, the ratio $\eta$$\pi^{0}$ reach a plateau value of 0.49, which is compatible with the measurement $0.457\pm 0.013\pm 0.018$ [57] and with the empirical fit result $0.487\pm 0.024$ [70].

### v3

@@ II Description of BWTPM
Figure 1: Fit of the midrapidity charged-particle multiplicity density data measured for various AA collisions at different centralities $x_{c}$ and energies $\sqrt{s_{\mathrm{NN}}}$ by the ALICE [17, 18, 19, 20], PHOBOS [21] and PHENIX [22] experiments. The dashed line shows the prediction (multiplied by 0.5 for better visibility) for the OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV. The data/fit ratio demonstrates the quality of the fits.
@@
To my knowledge, there is no generic parametrization in the literature similar to Eq. (13). The existing parametrizations of $M_{\rm ch}$ versus $x_{c}$ (or versus the number of participant nucleons $N_{\mathrm{part}}$) are valid only for a particular AA collision system [21, 30]. Of course, there are many sophisticated models describing the $M_{\rm ch}$ data, most of which use the hydrodynamical approach for particle production in AA collisions (see, e.g., Ref. [31]). However, such models are much more complex with respect to the simple parametrization of Eq. (13) and their accuracy is worse.
@@
Figure 2: Fit of the midrapidity charged-particle multiplicity density in inelastic pp collisions by Eq. (14) as a function of $\sqrt{s}$. Six higher energy data points are from Refs. [24, 25, 26] and the remaining data points are from Refs. [21, 27].
IV parametrization of BWTPM
@@
Here I present parametrizations for the model parameters versus $M_{\rm ch}$ (given by Eq. (13)), A, $\sqrt{s_{\mathrm{NN}}}$, and the particle mass and $p_{\textrm{T}}$. The parametrizations were chosen to be as simple as possible by examining the behavior of the model parameters versus the collision system, centrality, and energy, and using the results of Ref. [1] and other studies. For normalization parameters $V,V_{1}$, and $V_{2}$, to describe their strong increase with centrality [1] and taking into account the experimental fact that the hadron $p_{\textrm{T}}$-integrated yields increase linearly with $M_{\rm ch}$ at $M_{\rm ch}>100$ [32], the following parametrizations are assumed:
@@
Here and throughout the present paper the parameters $a_{i},i=1,2,3,$... are fitting constants. Parameter $f_{r}$ is nonzero for ${\mathrm{K}}^{*0}$and $\rho^{0}$ only, and accounts for the suppression of the yields of these short-lived resonances at low $p_{\textrm{T}}$ due to the rescatterings of their decay products in the hadronic medium. The suppression is stronger for more central collisions, and scales with $M_{\rm ch}^{1/3}$ [11].
The decrease of parameters $T$ and $k$ with increase of centrality and of parameters $e_{2}$ and $n$ with increase of centrality and $\sqrt{s_{\mathrm{NN}}}$ [1] can be parametrized as follows:
@@
The factor $f_{S}$ is equal unity for ordinary hadrons and increases with decreasing centrality (or $M_{\rm ch}$) for other hadrons. It is greater for hadrons containing larger number of strange and charm (anti)quarks, which are denoted by $n_{s}$ and $n_{c}$, respectively. For example, $n_{s}=1$ for K and $\eta$  mesons, and $n_{s}=2$ for the $\phi$ meson. A simple parametrization for $f_{S}$ in AA collisions is found
@@
which describes well the enhancement features of all considered hadrons containing at least one strange or charm (anti)quark ($n_{s}=n_{c}=0$ for ordinary hadrons). The term proportional to the hadron baryon number $B$ is introduced to better describe the production of hyperons. The same parameter $a_{21}$ for strange and charmed baryons is assumed since the data used for the latter case are very limited. For the hadrons containing bottom quarks, not considered in the present study, Eq. (20) can be modified accordingly. Note that $f_{S}$ tends to unity at $p_{\textrm{T}}>a_{22}$. Hence, the ratios of different particle yields in the present model are independent of centrality at high $p_{\textrm{T}}$, as in the data [11].
It is known that the $p_{\textrm{T}}$-averaged $v_{2}$ increases when going from central to peripheral collisions, and then decreases for very peripheral collisions [11, 16]. Also, the eccentricity $\varepsilon_{2}$ of the transverse geometry of the hadronic medium behaves similarly versus collision centrality. When going from central to peripheral collisions, this geometry goes from circular to elliptical and $\varepsilon_{2}$ increases. But the geometry goes again to circular and $\varepsilon_{2}$ decreases at very peripheral collisions, where on average only one nucleon-nucleon collision accurs [10, 15]. So one can assume $v_{2}\sim\varepsilon_{2}$. This parametrization of $r_{xy}$ ensures such a centrality dependence for $\varepsilon_{2}$ (given by Eq. (7))
@@
where $x=\textrm{A}^{a_{25}}(z_{c}+y)/(1+y)$, A is the atomic mass number, $z_{c}=M_{\rm ch}/M_{\rm ch}^{0\textrm{-}5\%}$ is the normalized multiplicity over the multiplicity of the 0-5% most central collisions $M_{\rm ch}^{0\textrm{-}5\%}$ (determined by Eq. (13) at $x_{c}=0.025$) and its correction $y=(a_{31}/\sqrt{s_{\mathrm{NN}}})^{a_{32}}z_{c}$ is needed mainly for the fit of $v_{2}$ data at RHIC energies. Parameter $f_{d}$ is introduced to describe a slightly different centrality dependence of $v_{2}$ data in Xe–Xe collisions compared to all other collision systems considered in present study for which $f_{d}=0$ is used. The difference is related to the non-spherical (deformed) shape of the 129Xe nucleus [33].
To describe the centrality and energy dependence of the hadrons $p_{\textrm{T}}$-spectra and also of the charged-particle average transverse momentum $\langle p_{\textrm{T}}\rangle$ [34, 35], particularly its local minimum at $z_{c}\approx 1.05$ and rise at $z_{c}>1.05$ for ultracentral collisions [35], I parametrized $\rho_{0}$ as
@@
Factor $f_{uc}$ [35, 36] is close to unity at $z_{c}\!\leq\!a_{38}\!\approx\!1.1$ and decreases at $z_{c}\!>\!a_{39}$. So, it increases $\rho_{0}$ at $z_{c}\!>\!a_{39}$ and therefore increases $\langle p_{\textrm{T}}\rangle$ for ultracentral events as in Refs. [35, 36]. Parameter $a_{35}$ is used to describe the observed slightly different centrality dependence of $\langle p_{\textrm{T}}\rangle$ in Xe–Xe and Pb–Pb collisions, which is due to the non-sphericity of 129Xe nuclei [34].
So, the parametrizations in Eqs. (16)–(IV) allow us to determine the aforementioned (in Sec. II) more than 600 independent values for $V,V_{1},V_{2},T,k,e_{2},n,f_{S},r_{xy},$ and $\rho_{0}$ by means of about 40 constants.
I also parametrized other ingredients of BWTPM. The energy dependences of parameter $f_{hyp}$ from Eqs. (11) and (12) and of the global normalization $f_{N}$ are given as
@@
where $a_{N}=1$ is assumed for Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}=$ 5.02 TeV, and for other used data sets the values of $a_{N}$ are defined relative to this one. As mentioned in Sec. II, it is assumed that the parameter $\rho_{2}$ as well as the functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ are proportional to $\varepsilon_{2}$ in order to ensure the relation $v_{2}\sim\varepsilon_{2}$ [13, 14, 15, 16]. A weak energy dependence of $\rho_{2}$ is parametrized as
$\rho_{2}=\varepsilon_{2}a_{44}/\sqrt{s_{\mathrm{NN}}}^{a_{45}}.$ (24)
Functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ are responsible mainly for the description of the high-$p_{\textrm{T}}$ behavior of elliptic flow $v_{2}$. For their dependence on $p_{\textrm{T}}$, $\sqrt{s_{\mathrm{NN}}}$, and $M_{\rm ch}$ I use
@@
The parametrization for mesons is inaccurate for $\mathrm{J}/\psi$.
@@
where $m_{\mathrm{p}}$ is the proton mass and the step function $H$ is defined as $H(x)=0(1)$ for $x\leq(>)~0$. So $c_{2}=1$ for pions and (anti)protons. Larger $a_{59}$ with respect to $a_{58}$ is nedded to describe the baryon yields. Note that the different parametrizations for mesons and baryons in Eqs.  (26) and (27) indicate the difference in production mechanisms of these hadrons containing two and three (anti)quarks, respectively.
@@
where parameter $\tilde{e}_{1}$ depends on the particle species. Parameter $f_{\mathrm{p}}$ is nonzero for the (anti)proton only to better describe its $p_{\textrm{T}}$-spectra at RHIC energies. Additional normalization parameter $f_{\rm D}$ is used for $\mathrm{D}^{\pm}$ mesons to describe their lower yield with respect to the ${\mathrm{D}}^{0}$ meson.
@@ V Fitted data and parameters
Thus, all the ingredients of the BWTPM are presented in previous sections. To determine the values of its free parameters, a global simultaneous fit is performed of the combined data on various hadrons midrapidity $p_{\textrm{T}}$-spectra, $\langle p_{\textrm{T}}\rangle$, and $v_{2}$ for different centrality classes in Pb–Pb collisions at energies $\sqrt{s_{\mathrm{NN}}}=2.76$ and 5.02 TeV [18, 34, 35, 37, 38, 41, 42, 43, 39, 40, 44, 45, 46, 49, 50, 47, 48, 51, 52, 55, 56, 53, 54, 57, 58, 59, 60, 61, 62] and in Xe–Xe collisions at 5.44 TeV [33, 63, 64, 65]. To check the validity of BWTPM for other colliding nuclei and at much lower energies, this fit also includes the $p_{\textrm{T}}$-spectra and $v_{2}$ for $\pi^{\pm}$, $\mathrm{K}^{\pm}$,  and $\mathrm{p}$($\mathrm{\overline{p}}$) measured in Au–Au collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV [66, 67], and the $p_{\textrm{T}}$-spectra for $\pi^{\pm}$ and $\mathrm{K}^{\pm}$ measured in Au–Au collisions at 39 and 27 GeV [68], in UU collisions at 193 GeV [69], and in Cu–Cu collisions at 200 GeV [70]. In the last four cases the proton data not included since they are not corrected for the contribution from weak decays of hyperons. The resulting ratio $\chi^{2}/NDF$ for this global fit (in the ROOT framework [28]) is 6738/13408. Values for the fitting parameters are given in Tables I–III. Parameters $f_{BW},f_{1}$, and $f_{2}$ for $\pi^{\pm}$ and $\pi^{0}$ are fixed equal to unity. Note that the fit uses the data of prompt $\rm D$} and $\mathrm{J}/\psi$ mesons and $\Lambda_{c}$ baryons, not including contributions from decays of heavier hadrons, containing $b$-quarks. To describe the $p_{\textrm{T}}$-spectra in Xe–Xe collisions [63, 64], the right-hand side of Eq. (12) is scaled by a factor of 1.07 to account for a small normalization difference between the measured $p_{\textrm{T}}$-spectra of identified [63] and unidentified charged particles [64]. Note also that parameters $a_{24}$$a_{32}$ and $a_{44}$$a_{57}$ are resposible mainly for the elliptic flow. Only some of them entering in Eq. (24) influence the azimuthally integrated $p_{\textrm{T}}$-spectra due to the $\rho_{2}$ contribution into the radial flow velocity Eq. (5). This contribution is small and decreases with increasing energy, according to Eq. (24). So, its influence on $p_{\textrm{T}}$-spectra is negligible at high energies and becomes sizable (at most about several % at high $p_{\textrm{T}}$) only for Au–Au collisions at low energies $\sqrt{s_{\mathrm{NN}}}=27$ and 39 GeV. It should be noted that all the parameters listed
@@
in Tables I–III have reasonable fitting uncertainties and are important for describing the data. For instance, the small parameter $a_{26}=0.0042\pm 0.0003$ is necessary for fitting $v_{2}$ for peripheral collisions. The global fit with $a_{26}=0$ has $\chi^{2}=7140$ and significantly overestimates the $v_{2}$ data (shown in Fig. 14) at $M_{\rm ch}<40$ and $x_{c}>0.7$.
@@ VI.1 Hadrons $p_{\textrm{T}}$-spectra, integrated yields, and $\langle p_{\textrm{T}}\rangle$
First, I present several plots to illustrate the results and quality of the $p_{\textrm{T}}$-spectra fits. Figure 3 displays the fits of different data sets with highest $p_{\textrm{T}}$ reached in central Pb–Pb collisions. To demonstrate the quality of the fits, the data points have been divided by the corresponding fit function values, and these ratios are also plotted in the bottom panel. Generally, the quality is good within the data uncertainties. Figures 4, 5, and 6 display the main fits of the centrality dependent spectra for the most abundantly produced pions, kaons, and protons. The fit
Figure 3: Fit of $p_{\textrm{T}}$-spectra with highest $p_{\textrm{T}}$ reached in Pb–Pb collisions for charged particles at $\sqrt{s_{\mathrm{NN}}}$ ==2.76 TeV [38, 41] and 5.02 TeV [38, 42] and for ${\mathrm{D}}^{0}$  at $\sqrt{s_{\mathrm{NN}}}$ ==5.02 TeV [43, 48].
Figure 4: Fit of pion, kaon, and proton $p_{\textrm{T}}$-spectra at $|y|<$ 0.5 for different centrality classes (centrality intervals) in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV [18]. The data points and fitting curves in the top panels are scaled, for better visibility, by the numbers given in the parentheses.
Figure 5: Similar to the Fig. 4 but for Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [63]. In addition, the dashed lines in the top panels show the predictions (scaled down by 0.0002 for better visibility) for 0-5 % centrality OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
@@
quality is almost always very good. This is true also for the fits of $p_{\textrm{T}}$-spectra of all considered other particle species. The predictions for the 0-5% most central OO collisions (using $a_{N}=1$) are shown in Fig. 5. Note that Figs. 3 and 4 repeat the similar ones of Ref. [1] to show that the present version of BWTPM with much fewer fitting parameters gives similar fit quality. Also, it is checked that the relative contributions of the BW, Ts, and Po components of the present model to the $p_{\textrm{T}}$-spectra are similar to the ones given in Ref. [1]. As shown in Table I, the contribution of the Ts-component is nonzero only for pions, kaons, and protons, which have significant feedback from the resonance decays.
@@
Then, Fig. 7 shows a good agreement of the model calculations (using the parameter values given in Sec. V) for the ratios of $p_{\textrm{T}}$-spectra of different hadrons with the ALICE data obtained in central and peripheral Pb–Pb collisions. These ratios, like the other ones given in Ref. [1], demonstrate more clearly the effect of hadron production mechanisms on the $p_{\textrm{T}}$-spectra. The peaks at $p_{\textrm{T}}$ $\approx$ 2–7 GeV/c are related to the radial flow, which is stronger for heavier particles and more central collisions. The ratios with the short-lived ${\mathrm{K}}^{*0}$ and $\rho^{0}$ resonances show the suppression of their yields for low $p_{\textrm{T}}$ and central collisions, according to Eq. (16). To demonstrate the collision system and energy dependence of particle ratios, the left panels of Fig. 7 show the model calculations for 0-10% centrality Au–Au (dashed lines) and Cu–Cu (dotted lines) collisions at $\sqrt{s_{\mathrm{NN}}}$ == 200 GeV. Both dependencies are not strong and appear at moderate $p_{\textrm{T}}$. Note that the hadron $p_{\textrm{T}}$-spectra in Eq. (II) have a power-law behavior $\propto p_{\textrm{T}}^{2-n}$ at high $p_{\textrm{T}}$. Hence, the ratios for all hadrons, which have $h_{n}=1$ in Table I, reach a plateau (independent of the collision system, centrality, and $\sqrt{s_{\mathrm{NN}}}$) at high $p_{\textrm{T}}$ where parameter $n$ in Eq. (19) depends only on $\sqrt{s_{\mathrm{NN}}}$. The plateau is simply defined by the ratio of factors $(2J+1)f_{2}$ with $f_{2}$ given in Table I. It is true also for the ratios between hyperons that have the same $h_{n}$. For example, the ratio $\eta$$\pi^{0}$ reach a plateau value of 0.49, which is compatible with the measurement $0.457\pm 0.013\pm 0.018$ [58] and with the empirical fit result $0.487\pm 0.024$ [71].